Har Ltd. is a logistics company operating in Subang Jaya. They have recently bid for a Selangor wide contract of providing medicines to all the health units. This is an excellent opportunity for the company to expand their business. To fulfill this requirement, Omair, who is the CEO of the company is analyzing, how he should expand his fleet. The basic question in front of him is whether he should buy his own fleet of trucks, or lease them. And also the question, whether he should buy them by raising capital through bonds or equities. He needs 100 trucks to fulfill the requirements. He has just finished a call with ORIX Leasing Ltd, who have said, that they would be willing to lease them 100 trucks for MYR 450,000 for 5 years. Instead, if Omair decides to buy the vehicles he is estimating that he would be spending MYR 18000 per truck. And the question is how to finance. He was checking the debt market rates, and found out that for similar risk companies, the bonds with a face value of 1000 were trading at MYR 940 with 5 years to maturity (annual coupon payment of 5%). Buying the trucks also means that he would be incurring a cost of MYR 85,000 each year for the whole fleet. The company’s tax rate is 30%. The tax laws allow straight-line depreciation for 5 years. The cost of capital is the same as the cost of debt.

Questions 1. What do you think whether Omair should BUY the trucks or LEASE them? Why?

2. Would it be wiser for him to raise equity instead of bonds IF he goes for Buy option?
Answer #1

Solution 1)

Option 1: Buy the trucks

Cost price of 1 truck = MYR 18,000

Cost price of 100 trucks at t=0 = MYR 18,000*100 = MYR 1,800,000

Depreciation on trucks = 1,800,000/5 = MYR 360,000

Maintenance cost per year = MYR 85,000

With the increase in cost, the company will be benefitted by lower profit before tax and would lead to lower tax. Hence, there would be tax benefit of 85000*(30%) = 25,500

Hence, net effective cash flow for maintenance = 85000 - 25500 = 59,500

Face value of the bond = MYR 1000

Price of the bond = MYR 940

Maturity = 5 years

Annual coupon rate = 5%

Annual coupon payment = 5% *1000 = MYR 50

Yield-to-Maturity.(YTM) can be calculated using the Rate function in Excel = Rate(Nper, Pmt, PV, FV,Type)

= Rate(5,50,-940,1000,0)

= 6.44%

Tax rate = 30%

Cost of debt = Yield-to-maturity (YTM) of debt *(1 - Tax rate%)

Cost of debt = 6.44%*(1 - 30%) = 6.44%*70% = 4.509%

Depreciation Tax Shield each year = Depreciation amount*Tax rate

= 360000*30% = 108,000

This is the benefit which the investor will get by buying the trucks

The cash flows are as follows:

Particulars t=0 t=1 t=2 t=3 t=4 t=5
After-tax Maintenance Costs -59500 -59500 -59500 -59500 -59500
Depreciation Tax Shield 108000 108000 108000 108000 108000
Net Operating Cash Flow 48500 48500 48500 48500 48500
Cost of the trucks -1800000
Net cash flow -1800000 48500 48500 48500 48500 48500
Net Present Value =NPV(Rate, Cash Flows from t=1) + Initial Investment
Net Present Value -1587139.94

Since NPV is negative, thus, Present value of outflow = 1587139.94

Option 2) Lease the trucks for MYR 450,000

Since lease rental is operating expense, hence, it would lead to lower profits before tax and lower tax expense

Thus, the after-tax rental cost at t=0 is = 450,000*(1 - 30%) = MYR 315,000

Since the present value of the rental is less, thus, Omair should go for lease option

Solution 2) Since the cost of equity is more than the cost of debt. At lower cost also, the better option is to go for the lease, hence, if Omair takes the equity then cost will increase which would make the NPV more negative.   

Please comment in case of any doubts or clarifications required. Please Thumbs UP!!